OpenStudy (anonymous):

What will be length of AE? 3 cm 4 cm 5cm 7 cm

OpenStudy (anonymous):

diagram?

OpenStudy (anonymous):

5

OpenStudy (anonymous):

3 centimeters.

OpenStudy (anonymous):

14/7 = 10/AE 2 = 10/AE AE = 10/2 = 5

OpenStudy (anonymous):

how did you got 3 sm @nightshifting

OpenStudy (anonymous):

My bad, KML is right xD

OpenStudy (anonymous):

Sorry about that.

OpenStudy (anonymous):

np it's ok!

OpenStudy (anonymous):

you got it @Heatherrr

OpenStudy (anonymous):

Kml, is right @Heatherrr. Best answer.

OpenStudy (anonymous):

thanks yall

OpenStudy (anonymous):

uw